OpenStudy (anonymous):

what is the formula for the distance between skew planes? Is it the same formula as the distance between parallel planes?

OpenStudy (lgbasallote):

skew planes?

OpenStudy (lgbasallote):

did you type that right? it's really planes?

OpenStudy (anonymous):

Yes Quoting from the Stewart Calculus textbook, "lines L1: x = 1 +t y = -2 +3t z =4 -t L2: x =2s y = s z =-3 +4s are skew. Find the distance between them. Solution "Since L1 & L2 are skew, they can be viewed as lying on 2 parallel planes P1 & P2. The distance between L1 & L2 is the same as the distance between P1 & P2". They say it can be computed as D = abs ax1 +by1 +cz1 +d/ squrt (a^2 +b^2 + c^2)

OpenStudy (anonymous):

The planes are not skew, the lines are. Perhaps you are having a little trouble visualizing 2 parallel planes containing them.

OpenStudy (anonymous):

thanks. Yes I do have trouble visualising.  However, I have found a drawing that illustrates what you said at this link. http://intermath.coe.uga.edu/dictnary/descript.asp?termID=424